# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ 3.3 Identifiability of next-token predictors
Now assume that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta_{d}$ and $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta_{\tilde{d}}$ are both maximizers of 3. If both models have enough capacity to represent the ground-truth data distribution $p_{\mathcal{D}}$, then they necessarily represent the same distribution, i.e. $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\mathcal{D}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. By Theorem 5, we can thus conclude that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{{\bm{\mathrm{g}}}})$. Assuming these assumptions hold in practice, this would imply that all models trained with sufficient capacity on a given data distribution $p_{\mathcal{D}}$ will be $\sim_{EL}$-equivalent. This analysis relies crucially on the assumption that $\Theta$ has enough capacity to represent $p_{\mathcal{D}}$. In practice, this might not hold for at least two reasons. First, we typically train with a fixed representation dimension $d$ which limits the expressivity of the model. Moreover, despite the universal approximation guarantee, there might not even exist a sufficiently large $d$ such that the model can express $p_{\mathcal{D}}$ exactly. Secondly, all distributions that can be represented by $\Theta$ put nonzero probability mass on all text sequences (because of the $\exp$), whereas one can argue that the ground-truth distribution of text $p_{\mathcal{D}}$ gives zero probability to many sequences.
@@ 4 Linear properties
We will now focus on linear properties of language models. These describe how a given model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ represents different inputs—similar to our opening example, illustrating a geometric relationship (parallelism) among the vector differences between the embeddings of two different inputs (“lucky” and “luckiest”) and that between two further inputs (“easy” and “easiest”). These linear properties are not to be confused with the linear equivalence class $\sim_{L}$ mentioned in Corollary 6, which instead describes how different models represent the same data distribution. In Section 5, we will combine the linear properties defined here with our equivalence class of next-token predictors, described in Section 3, to study what properties are shared by all models within the same equivalence class. See also Figure 1 for an illustration.
Context-query-reply sequences.
To describe linear properties, we use a set up similar to Paccanaro and Hinton [1], Hernandez et al. [2]. We consider sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ that can be decomposed as ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$, where ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ denotes the context (or subject), ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ denotes the query (or relation), and $y\in\mathcal{A}$ is the reply (or object). The following example showcases the setting.
Example 2.
We consider sentences of the form ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$ where, for example, ${\bm{\mathrm{s}}}=$“All roads lead to Rome”, ${\bm{\mathrm{q}}}=$“What is the written language?”, with most likely prediction $y=$“English”. This resembles a setup where the model is tested to determine whether it understands the property of “written language”. Another example could be ${\bm{\mathrm{s}}}=$“Rome”, ${\bm{\mathrm{q}}}=$“is the capital of” and $y=$“Italy”.
Sentences ${\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$ like these can be found in natural language, capturing relational aspects between ${\bm{\mathrm{s}}}$ and $y$, mediated by ${\bm{\mathrm{q}}}$. In our setting, though, formally these tuples allow for arbitrary ${\bm{\mathrm{s}}}$ and ${\bm{\mathrm{q}}}$, including queries whose expected reply is independent of the context, like ${\bm{\mathrm{q}}}=$Whatever follows reply with $42$.”, or paraphrases, e.g., ${\bm{\mathrm{q}}}^{\prime}=$Now reply with $42$!”. In Appendix E, we demonstrate how our analsys meaningfully captures these corner cases.
4.1 Relational linear property
We describe the linear structure of representations solely relationally, that is, in terms of context-query-reply sequences ${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$, similar to Paccanaro and Hinton [1]. Relatedly, Hernandez et al. [2] provided empirical evidence for relational linearity between hidden transformer layers.
Our analysis focuses on embeddings ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})\in\mathrm{SIm}({\bm{\mathrm{f}}})$ and subspaces $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, allowing us to define relational linear properties solely in terms of the quantities described in our identifiability result; this enables our analysis to be agnostic to assumptions on the data generating process underlying natural language and, e.g., to not require positing nor referring to unobserved variables.
For a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$, let $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ be a subspace. We say that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents the query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ on $\Gamma$, if, there exist a matrix ${\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d\times d}$ and a vector ${\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d}$ such that, for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$,
When this holds, we indicate with $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$.
Intuitively, all the information relevant within ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ to compute the probability of tokens in $\Gamma$ is captured, up to an affine transformation, by ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ in the subspace $\Gamma_{\bm{\mathrm{q}}}$. Indeed, one can show that, if ${\bm{\mathrm{g}}}_{0}({\bm{\mathrm{y}}})\in\Gamma$, then it must be that ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)=\big{(}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\big{)}^{\top}{\bm{\mathrm{g}}}_{0}(y)$. The space $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$ are central to proving whether relational linearity holds for all $\sim_{EL}$-equivalent models, as we will show in Section 5.
4.2 Connection to other linear properties
Here, we show how to specialize the notion of relational linearity to capture three linear properties that have been empirically observed. We follow the taxonomy by Park et al. [23]. All proofs are reported in Appendix C.
4.2.1 Linear subspaces (ls)
@@
We show next that parallel vectors induce similar log ratios of conditional probabilities, as noted in [23, 28]:
@@
Take a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. For $y_{0},y_{1},y_{2},y_{3}\in\mathcal{A}$, the difference vectors ${\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0})$ and ${\bm{\mathrm{g}}}(y_{3})-{\bm{\mathrm{g}}}(y_{2})$ are parallel in $\mathcal{N}$ if and only if there exists $\beta\neq 0,\text{ s.t. }\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ it holds that
@@
That is, parallel difference vectors for token pairs correspond to proportional likelihood ratios between the tokens in each token pair. Parallel vectors can be associated with their one-dimensional subspace, and are possibly related to changes in replies for a query ${\bm{\mathrm{q}}}$. We can capture this through the following definition:
Definition 10 (ls: Relational Linear Subspaces).
A model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ linearly represents a subspace ${\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$ relative to ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ if, for all couples of tokens $y_{i},y_{j}\in\mathcal{A}$ such that ${\bm{\mathrm{g}}}_{i}(y_{j}):={\bm{\mathrm{g}}}(y_{j})-{\bm{\mathrm{g}}}(y_{i})\in\Gamma$, there exists a vector ${\bm{\mathrm{\gamma}}}\in\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ such that $\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ it holds that:
An example is visualized in Figure 3. We show that the ls property is implied by relational linearity in the following sense:
[…]
Suppose that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (1) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, and also (2) ${\Gamma_{\bm{\mathrm{q}}}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$, then the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents $\Gamma_{\bm{\mathrm{q}}}$ relative to ${\bm{\mathrm{q}}}$ (Definition 10).
@@ 4.2.2 Linear probing (lp)
There is empirical evidence for sentence embeddings being linearly separable into embeddings of text of different languages [23, 29]. Such a structure also enables what is called linear probing [30, 31] (see below). We define linear probing as relational property:
@@
We say that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ can be linearly probed for a query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ and a collection $\mathcal{Y}_{P}\subseteq\mathcal{A}$ of $\ell$ elements if there exist ${\bm{\mathrm{W}}}\in\mathbb{R}^{\ell\times d}$ and ${\bm{\mathrm{b}}}\in\mathbb{R}^{\ell}$, such that for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ and $\forall i\in[\ell]$ it holds that
@@
where $p(y\mid\cdot\,;\mathcal{Y}_{P})=p(y\mid\cdot\,)/(\sum_{y^{\prime}\in\mathcal{Y}_{P}}p(y^{\prime}\mid\cdot\,))$ denotes the conditionals restricted to $\mathcal{Y}_{P}$.
To illustrate why this is termed linear probing, suppose a model given the query ${\bm{\mathrm{q}}}=$Is the text written in English?” discriminates input sequences ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ between positive $y_{0}=$yes” and negative examples $y_{1}=$no”, that is it assigns high probability mass to either $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ or to $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$. Then, these conditional distributions can be evaluated directly from ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ via a linear probe. See Figure 1 for an illustration. We show that $\Gamma$lr allows for lp:
@@
If a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (1) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, and (2) ${\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j})\in\Gamma$ for all $y_{i}\in\mathcal{Y}_{P}$, then the model can be linear probed (Definition 12) for ${\bm{\mathrm{q}}}$ and $\mathcal{Y}_{P}$, with parameters given by ${\bm{\mathrm{W}}}=({\bm{\mathrm{w}}}_{1},\ldots,{\bm{\mathrm{w}}}_{\ell})^{\top}$ and ${\bm{\mathrm{b}}}=(b_{1},\ldots,b_{\ell})^{\top}$, where ${\bm{\mathrm{w}}}_{i}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and $b_{i}:=({\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}(y_{i})$.

### v2

@@ 3.4 Implications for empirical practice
Suppose that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta_{d}$ and $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta_{\tilde{d}}$ are both global maximizers (in their respective model classes $\Theta_{d}$ and $\Theta_{\tilde{d}}$) of the objective in Equation (3). If both models have enough capacity to represent the ground-truth data distribution $p_{\mathcal{D}}$, then they necessarily represent the same distribution, i.e., $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\mathcal{D}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. By Theorem 5, we can thus conclude that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{{\bm{\mathrm{g}}}})$. If these assumptions held in practice, this would imply that all models trained with sufficient capacity on a given data distribution $p_{\mathcal{D}}$ will be $\sim_{EL}$-equivalent. [^2] We neglect optimiziation issues such as local minima for ease of exposition. This analysis relies crucially on the assumption that $\Theta$ has enough capacity to represent $p_{\mathcal{D}}$. In practice, this might not hold for at least two reasons. First, we typically train with a fixed representation dimension $d$ which limits the expressivity of the model; moreover, despite the universal approximation guarantee, there might not even exist a sufficiently large $d$ such that the model can express $p_{\mathcal{D}}$ exactly. Secondly, all distributions that can be represented by $\Theta$ put nonzero probability mass on all text sequences (because of the exponential on the RHS of Equation (1)), whereas under the ground-truth distribution $p_{\mathcal{D}}$ describing, e.g., text on the internet, several sequences will have zero probability. Models of the form in Equation (1) are thus inherently misspecified in such cases. Modeling these settings might thus require an extension of the current theoretical framework [Nielsen et al., 2024].
Different token vocabularies.
Our analysis is also restricted to next-token predictors that share the same token vocabulary $\mathcal{A}$. We hypothesize that, for two models with different token vocabularies $\mathcal{A}$ and ~$\tilde{}\mathcal{A}$, our results may be extended to prove a $\sim_{EL}$-equivalence relation restricted to the shared tokens $\mathcal{A}\cap\tilde{}\mathcal{A}$, under suitable conditions on the next-token probabilities.
@@ 4 Linear properties
In the previous section, we established identifiability results for next-token predictors. Here, we turn to precisely defining the linear properties we will focus on. These characterize how a given model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ represents different inputs—as in our opening example, describing a geometric relationship (parallelism) among the vector differences between the embeddings of two different inputs (“lucky” and “luckiest”) and that between two further inputs (“easy” and “easiest”). Importantly, these linear properties are not to be confused with the linear equivalence class $\sim_{L}$ in Definition 3, which instead describes how different models represent the same data distribution. In Section 5, we will combine the linear properties defined here with the identifiability results of Section 3 to determine which linear properties hold for all models in a given equivalence class. See also Figure 1 for an illustration.
Our analysis focuses on embeddings ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})\in\mathrm{SIm}({\bm{\mathrm{f}}})$ and unembeddings ${\bm{\mathrm{g}}}_{0}(y)\in\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, allowing us to define relational linear properties solely in terms of the quantities described in our identifiability result: our analysis is thus agnostic to assumptions on the data-generating process underlying natural language and it does not require positing unobserved variables. In principle, the linear properties we will define can apply to any collection of strings—for example, the difference between the unembeddings of “1fv0sywi” and “eg2op3te” could be parallel to the difference between those of “tgsqil2h” and “khdo5zof”. As we will argue, such parallel structures imply certain symmetries in a model’s conditional next-token probabilities. Our work is motivated by the commonly observed instances of linear properties involving collections of semantically meaningful strings, where symmetries in next-token probabilities likely reflect regularities in human-produced text.
4.1 Parallel vectors
@@
We next show that parallel vectors induce similar log ratios of conditional probabilities, as noted in [Park et al., 2024a, Jiang et al., 2024]:
@@
Consider a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. For $y_{0},y_{1},y_{2},y_{3}\in\mathcal{A}$, the difference vectors ${\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0})$ and ${\bm{\mathrm{g}}}(y_{3})-{\bm{\mathrm{g}}}(y_{2})$ are parallel in $\mathcal{N}$ if and only if there exists $\beta\neq 0,\text{ s.t. }\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$
@@
That is, parallel difference vectors for token pairs correspond to proportional likelihood ratios between the tokens in each token pair. Notice that, as in Definition 7, the difference vectors are parallel only in the space $\mathcal{N}$. This implies that the components outside $\mathcal{N}$ for two $\mathcal{N}$-parallel vectors are allowed to not be parallel. All proofs for the results presented in this section are provided in Appendix C.
@@
4.2 Relational linear property
Beyond parallelism, the first property we define is relational linearity, introduced by Paccanaro and Hinton [2001] and recently studied by Hernandez et al. [2024], who found empirical evidence that hidden transformer layers in language models display this property.
Context-query-reply sequences.
We consider sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ that can be decomposed as ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$, where ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ is termed context (or subject), ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ is termed query (or relation), and $y\in\mathcal{A}$ is termed reply (or object). The following example illustrates a semantically meaningful context-query-reply sequence.
Example 2.
Consider a sequence ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$ where ${\bm{\mathrm{s}}}=$“All roads lead to Rome”, ${\bm{\mathrm{q}}}=$“What is the written language?”, and $y=$“English”. We deliberately pick $y=$“English” as the most likely next-token prediction following ${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}$ made by English speakers. The string ${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}$ could also be provided as input to a language model to test whether it can recognize English language. Another example of context-query-reply sequence is ${\bm{\mathrm{s}}}=$“Rome”, ${\bm{\mathrm{q}}}=$“is the capital of” and $y=$“Italy”.
[…]
For a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$, let $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ be a subspace. We say that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents the query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ on $\Gamma$, if there exist a matrix ${\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d\times d}$ and a vector ${\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d}$ such that, for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$,
When this holds, we define $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$.
Intuitively, all the information within ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ which is relevant to compute the probability of next-tokens in $\Gamma$ is captured, up to an affine transformation, by ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ in the subspace $\Gamma_{\bm{\mathrm{q}}}$. Indeed, one can show that, if ${\bm{\mathrm{g}}}_{0}({\bm{\mathrm{y}}})\in\Gamma$, then necessarily ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)=({\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$. The spaces $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$ are central to proving whether relational linearity holds for all $\sim_{EL}$-equivalent models, as we will show in Section 5.
Connection to other linear properties. In the following, we show how to capture three additional linear properties building on the definition of relational linearity. We follow the taxonomy by Park et al. [2024a].
4.2.1 Linear subspaces (ls)
[…]
Definition 10 (ls: Relational Linear Subspaces).
A model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ linearly represents a subspace ${\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$ relative to ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ if for all pairs of tokens $y_{i},y_{j}\in\mathcal{A}$ such that ${\bm{\mathrm{g}}}_{i}(y_{j}):={\bm{\mathrm{g}}}(y_{j})-{\bm{\mathrm{g}}}(y_{i})\in\Gamma$, there exists a vector ${\bm{\mathrm{\gamma}}}\in\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ such that ${\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})}$
We provide one example of this property in Figure 3. The ls property is implied by relational linearity (Definition 9) in the following sense:
[…]
Suppose that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (i) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, and (ii) ${\Gamma_{\bm{\mathrm{q}}}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$, then the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents $\Gamma_{\bm{\mathrm{q}}}$ relative to ${\bm{\mathrm{q}}}$ (Definition 10).
@@ 4.2.2 Linear probing (lp)
There is empirical evidence that, in language models, sentence embeddings can be linearly separated with good accuracy based on the language of the corresponding sentences [Park et al., 2024a, Chang et al., 2022].This property is also termed linear probing [Alain, 2016, Kim et al., 2018]. Below, we redefine linear probing as a relational property, based on Definition 9:
@@
We say that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ can be linearly probed for a query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ and a collection $\mathcal{Y}_{P}\subseteq\mathcal{A}$ of $\ell$ elements if there exist ${\bm{\mathrm{W}}}\in\mathbb{R}^{\ell\times d}$ and ${\bm{\mathrm{b}}}\in\mathbb{R}^{\ell}$ such that for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ and $\forall i\in[\ell]$
@@
where $p(y\mid\cdot\,;\mathcal{Y}_{P})=p(y\mid\cdot\,)/(\sum_{y^{\prime}\in\mathcal{Y}_{P}}p(y^{\prime}\mid\cdot\,))$ is the conditional distribution restricted to the set $\mathcal{Y}_{P}$.
To illustrate why this is termed linear probing, suppose a model given the query ${\bm{\mathrm{q}}}=$Is the text written in English?” discriminates input sequences ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ between positive $y_{0}=$yes” and negative examples $y_{1}=$no”—that is, it assigns high probability ${p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ to sequences ${\bm{\mathrm{s}}}$ corresponding to English sentences, and high probability $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ to non-English sentences. Then, these conditional distributions can be evaluated directly from ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ via a linear probe. Figure 1 includes an illustration of lp. Below, we relate lp (Definition 12) to $\Gamma$lr (Definition 9):
@@
If a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (i) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, and (ii) ${\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j})\in\Gamma$ for all $y_{i}\in\mathcal{Y}_{P}$, then the model can be linear probed (Definition 12) for ${\bm{\mathrm{q}}}$ and $\mathcal{Y}_{P}$, with parameters given by ${\bm{\mathrm{W}}}=({\bm{\mathrm{w}}}_{1},\ldots,{\bm{\mathrm{w}}}_{\ell})^{\top}$ and ${\bm{\mathrm{b}}}=(b_{1},\ldots,b_{\ell})^{\top}$, where ${\bm{\mathrm{w}}}_{i}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and $b_{i}:=({\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}(y_{i})$.